2 INGRID

3 Masashi Otani 4 Akira Murakami 5 Akihiro Minamino

6 Kyoto University

7 October 20, 2010

8 Contents

9 1 Monte Carlo simulation 3 10 1.1 Overview about MC . . . . . . . . . . . . . . . . . . . . . . . . . 3 11 1.2 Check detector response MC with Beam data . . . . . . . . . . . 5 12 1.3 Efficiency to neutrino interaction . . . . . . . . . . . . . . . . . . 5 13 1.4 Expected number of neutrino observation . . . . . . . . . . . . . 5

14 Chapter 1

15 Monte Carlo simulation

16 In this chapter, we explain about Monte Carlo simulation (MC) currently used 17 in INGRID analysis.

18 • Overview about MC

19 • Check detector response MC with Beam Data.

20 • Efficiency to neutrino interaction.

21 • Expected # of neutrino observation

22 1.1 Overview about MC

23 INGRID MC is composed of three parts : Jnubeam, NEUT and Detector re-24 sponse (figure fig:mcoverview).

25 • Neutrino Flux : Jnubeam (version 10c)

26 • Neutrino interaction to Target : NEUT (version 5.0.6.)

27 • Detector response to generated particles from neutrino interaction : Sim-28 ulator based GEANT4 (Detector response MC)

29 This INGRID MC is not software of ND280 software packages (nd280mc). 30 Integration of INGRID MC to ND280 packages is on going now. we plan to use 31 nd280mc including INGRID in future.

32

33 Jnubeam is T2K neutrino-beam line simulation (based GEANT3). We make 34 neutrino flux (ntuple-based flux), which is called as Jnubeam flux, to INGRID 35 with this simulator. In current INGRID MC (2010.Oct), we use version 10c of 36 Jnubeam mainly.

Figure 1.1: INGRID MC overview

37 Then, we simulate the interaction between the each flavor of neutrino ob-38 tained from Jnubeam flux files and target nucleus with NEUT (use 5.0.6. ver-39 sion). This target of INGRID is Iron nucleus(Fe) or scintillator nucleus(CH), 40 but now we use the interaction to Iron mainly. About interaction to CH, we are 41 progress in mass-production and study. 42 Finally, we simulate detector response to generated particles from neutrino 43 interaction with simulator based GEANT4 which is developed by Japanese of 44 INGRID group. We obtain the neutrino interaction vertex-X and vertex-Y from 45 neutrino vertex of Jnubeam flux file (these variable names are ”xnu” and ”ynu”). 46 The vertex-Z is uniform in each module, but distribution of the vertex in Iron 47 and scintillator is weighted with the mass ratio of Iron planes (99.54 ton) to 48 scintillator planes (3.74 ton). The detector response MC does not cover the 49 whole detector response of INGRID perfectly, but includes some parts which 50 have an impact to the efficiency to neutrino interaction mainly. Including parts 51 is below,

52 • Quenting effect of scintillator and attenuation of photon propagating in 53 the fiber.

54 • MPPC response model (including the effect of cross-talk and after pulse, 55 and the effect of pixel saturation).

56 • Real geometry of scintillator bar (effect on tracking efficiency).

57 For MPPC response model, we refer to page 11 of the slide ”Characterisa-58 tion of MPPC linearity response with the TRIP-T electronics ” (reported by 59 Calibration group of ND280 working group in 2009. this slide put at t2k.org). 60 We tuned the scale of exchange from energy to photon of MC with beam 61 related sand muon. We set this scale to adjust the peak p.e. deposit by muon 62 generated in MC to the peak p.e. deposit by sand muon. This tuning of scale 63 factor is just temporary, so need more tuning this scale to refine the estimation 64 of photon generated at MC (but, in currently analysis, the p.e. threshold is not 65 so much critical to the efficiency to neutrino).

66 1.2 Check detector response MC with Beam data

67 We compare the detector response to neutrino of MC to one of beam data at 68 some basic distribution ( number of active planes, reconstructed vertex, recon-69 structed tracking angle). The beam data set is Run2010a(from January to June, 70 in 2010). Total number of protons used in analysis is 3.26e19 and total number 71 of good spills is 1005887. 72 The results of comparison is showed in other section.

73 1.3 Efficiency to neutrino interaction

74 We calculate the efficiency to the neutrino interacted Fiducial volume (FV) at 75 each mode (CC+NC, CC, CCQE, CC-nonQE, NC). The error of efficiency is 76 including only MC statistics error (figure1.2, figure1.3).

Figure 1.2: Efficiency to ν_{μ} interacted in FV. Black point show the efficiency to all interaction mode. Red points show one to only CC interaction mode. Blue point show one to only NC interaction mode. The error includes only MC statistics error.

Figure 1.3: Efficiency to ν_{μ} interacted in FV. Red points show the efficiency to CCQE interaction mode. Blue points show one to CC-nonQE interaction mode. The error includes only MC statistics error.

77 Now we are creating the efficiency plot to other neutrino flavor (ν̄_{μ}, ν_{e},ν̄_{e}) ( 78 mass-production of MC was done for ν_{μ}, ν̄_{μ}. About ν_{e},ν̄_{e}, to be prepared).

79 1.4 Expected number of neutrino observation

80 We estimate the expected number of neutrino observed at INGRID with this 81 MC.We estimate three value: integrated flux to each modules, the number of 82 interaction in each modules, the number of observation at each modules. And 83 we estimate with other primary hadron production model (FLUKA2008) for 84 collision between proton beam and target carbon at Jnubeam. The default 85 hadron production model of Jnubeam(10c) is GCALOR/GFLUKA. We check 86 the effect of the hadron production difference on the each number. We show 87 the number of integrated flux, interaction, observation at each module (from 88 table tab:fluxnumu to table tab:obsnumubar). We summarize the total number 89 of neutrino observed at INGRID to table tab:sumexp and compare the total 90 number of MC to one of Data (Run2010a). Now, we use the ν_{μ} and ν̄_{μ} to 91 calculate the expectation. About ν_{e} and ν̄_{e}, we are progress in mass-production. 92 ν_{e} and ν̄_{e} contribution to total observation is less than 1% at flux.

Table 1.1: Integrated ν_{μ} flux at each modules [×10^{17}/10^{21}P OT]
module  0  1  2  3  4  5  6
1.GCALOR/GFLUKA 5.21 6.18 6.87 7.12 6.88 6.18 5.12
2.FLUKA2008  4.84 5.64 6.19 6.39 6.17 5.60 4.82
Ratio (2./1.) [%]  92.8 91.2 90.0 89.7 89.7 90.6 92.5
module  7  8  9  10 11 12 13
1.GCALOR/GFLUKA 5.44 6.40 7.11 7.36 7.14 6.44 5.50
2.FLUKA2008  5.08 5.86 6.45 6.66 6.41 5.81 5.05
Ratio (2./1.) [%]  93.4 91.6 90.8 90.5 89.8 90.1 91.7

Table 1.2: Integrated ν̄_{μ} flux at each modules [×10^{17}/10^{21}P OT]

module  0  1  2  3  4  5  6
1.GCALOR/GFLUKA 0.342 0.395 0.436 0.455 0.438 0.395 0.342
2.FLUKA2008  0.234 0.254 0.274 0.284 0.274 0.254 0.235
Ratio (2./1.) [%]  68.4 64.1 62.7 62.3 62.7 64.4 68.7
module  7  8  9  10  11  12  13
1.GCALOR/GFLUKA 0.371 0.415 0.452 0.474 0.454 0.417 0.375
2.FLUKA2008  0.251 0.271 0.284 0.286 0.282 0.270 0.253
Ratio (2./1.) [%]  67.6 65.4 62.9 60.3 62.1 64.8 67.6

Table 1.3: number of ν_{μ} interacted at each modules [×10^{6}/10^{21}P OT]

module  0  1  2  3  4  5  6
1.GCALOR/GFLUKA 2.73 3.55 4.13 4.33 4.19 3.58 2.71
2.FLUKA2008  2.25 2.86 3.29 3.46 3.31 2.84 2.23
Ratio (2./1.) [%]  82.3 80.4 79.6 79.8 79.0 79.3 82.3
module  7  8  9  10 11 12 13
1.GCALOR/GFLUKA 2.89 3.73 4.29 4.48 4.32 3.76 2.93
2.FLUKA2008  2.40 3.01 3.45 3.62 3.43 2.97 2.37
Ratio (2./1.) [%]  83.0 80.7 80.5 80.9 79.5 79.1 80.6

Table 1.4: number of ν̄_{μ} interacted at each modules [×10^{4}/10^{21}P OT]

module  0  1  2  3  4  5  6
1.GCALOR/GFLUKA 8.36 10.9 13.0 14.2 13.1 10.4 7.95
2.FLUKA2008  3.89 4.55 5.44 5.92 5.42 4.55 3.79
Ratio (2./1.) [%]  46.5 41.7 41.9 41.6 41.2 43.7 47.7
module  7  8  9  10 11 12 13
1.GCALOR/GFLUKA 9.58 11.1 14.1 14.7 13.3 11.7 9.34
2.FLUKA2008  4.37 4.91 5.61 5.58 5.40 5.19 4.47
Ratio (2./1.) [%]  45.6 44.1 39.7 37.9 40.5 44.5 47.8

Table 1.5: number of ν_{μ} observed at each modules [×10^{6}/10^{21}P OT]

module  0  1  2  3  4  5  6
1.GCALOR/GFLUKA 0.863 1.19 1.43 1.48 1.45 1.21 0.870
2.FLUKA2008  0.665 0.901 1.08 1.12 1.08 0.898 0.671
Ratio (2./1.) [%]  77.1 75.5 75.2 75.6 74.6 73.9 77.1
module  7  8  9  10 11  12  13
1.GCALOR/GFLUKA 0.936 1.27 1.48 1.54 1.48 1.27 0.951
2.FLUKA2008  0.729 0.961 1.13 1.19 1.11 0.945 0.716
Ratio (2./1.) [%]  77.9 75.9 76.1 76.9 75.3 74.1 75.3

Table 1.6: number of ν̄_{μ} observed at each modules [×10^{4}/10^{21}P OT]
module  0  1  2  3  4  5  6
1.GCALOR/GFLUKA 3.09 4.09 4.93 5.23 4.91 3.88 2.87
2.FLUKA2008  1.38 1.63 1.99 2.11 1.96 1.61 1.35
Ratio (2./1.) [%]  44.7 39.9 40.4 40.4 39.8 41.5 46.9
module  7  8  9  10 11 12 13
1.GCALOR/GFLUKA 3.54 4.11 5.22 5.77 4.78 4.49 3.46
2.FLUKA2008  1.48 1.81 1.99 2.03 1.89 1.99 1.62
Ratio (2./1.) [%]  41.7 44.1 38.1 35.1 39.6 44.3 46.8

Table 1.7: Expected total number of neutrino observation at INGRID [/10^{14}P OT]

ν_{μ}  ν̄_{μ}  ν_{μ}
GCALOR/GFLUKA 1.72 0.0559 1.78
FLUKA2008  1.38 0.0263 1.41

Table 1.8: Comparison of total number of neutrino observation between Data and MC. This MC includes only ν_{μ} and ν̄_{μ} data. MC of ν_{e} and ν̄_{e} is to be prepared.

Observation [/10^{14}P OT] Ratio to Data
Data  1.52  1
GCALOR/GFLUKA  1.78  1.17
FLUKA2008  1.41  0.88